# Provably Secure Disambiguating Neural Linguistic Steganography

### v1

@@ Abstract
Recent research in provably secure neural linguistic steganography has overlooked a crucial aspect: the sender must detokenize stegotexts to avoid raising suspicion from the eavesdropper. The segmentation ambiguity problem, which arises when using language models based on subwords, leads to occasional decoding failures in all neural language steganography implementations based on these models. Current solutions to this issue involve altering the probability distribution of candidate words, rendering them incompatible with provably secure steganography. In this paper, we propose a novel secure disambiguating method named SyncPool, which effectively addresses the segmentation ambiguity problem. We first group all tokens with prefix relationships in the candidate pool before the steganographic embedding algorithm runs to eliminate uncertainty among ambiguous tokens. In order to enable the receiver to synchronize the sampling process of the sender, a shared cryptographically secure pseudorandom number generator (CSPRNG) is deployed to select a token from the ambiguity pool. SyncPool does not change the size of the candidate pool or the distribution of tokens within the pool and thus is applicable to all provably secure language steganography methods. We also provide theoretical proofs. We experimentally demonstrate the applicability of our solution to various languages and models, showing its potential to significantly improve the reliability and security of neural linguistic steganography systems.
@@ II-A Steganography System and Steganographic Security
$\left|\text{Pr}\left[\mathcal{A}_{\mathcal{D}}^{\textsc{Encode}_{\mathcal{D}}\left(K,\cdot,\cdot\right)}=1\right]-\text{Pr}\left[\mathcal{A}_{\mathcal{D}}^{\mathcal{O}_{\mathcal{D}}\left(\cdot,\cdot\right)}=1\right]\right|<\text{negl}\left(k\right),$ (4)
where $\mathcal{O}_{\mathcal{D}}\left(\cdot,\cdot\right)$ is an oracle that can randomly sample from the channel distribution $\mathcal{D}$, $k$ is the security parameter of the shared key $K$ (usually the length of $K$), and $\text{negl}\left(k\right)$ is a negligible function concerning $k$.
@@ II-C Segmentation Ambiguity
The commonly used high-performance Transformer-based language models, represented by the GPT series [23, 24, 37] and the LLaMA series [35, 38], mostly utilize subword tokenization [27, 28, 29] to model text. Subword tokenization is a technique in which a word is split into subwords, and these subwords are known as tokens. This technique is used because a generative language model needs to maintain a large vocabulary and complex word structures. The concept behind this is that frequently occurring words should be in the vocabulary whereas rare words are split into frequent subwords. Each token is assigned an ID as a numerical representation of the subword  [^1] Since the numeric IDs corresponding to tokens differ across different models, in this paper, we directly use the subword mapped by the token to represent the token.. For example, the word “_unwanted” might be split into “_un”, “want”, and “ed”. The stego generated by the steganographic encoding algorithm $\textsc{Encode}_{\mathcal{D}}$ is essentially a sequence composed of stego tokens. The sender must detokenize it using a tokenizer into a stegotext before transmission. The reason is obvious because even in today’s prevalent use of generative models and generated text, transmitting tokens directly over a public channel is a highly suspicious behavior. For example, if the sender generates two stego tokens mapping to subwords “_any” and “thing”, the sender needs to detokenize them into the text “anything” before sending it to Bob. However, the issue is that common words like “anything” often exist as independent tokens “_anything” in the model’s vocabulary as well. As a result, a single piece of text can correspond to two or even more different token representations. This phenomenon is referred to as segmentation ambiguity. To replicate Alice’s generation process and extract the message using $\textsc{Decode}_{\mathcal{D}}$, Bob must retokenize the received text into tokens. But since both “_any” and “_anything” exist in the candidate pool, Bob cannot determine which token the sender embedded the message into. Clearly, the differing tokenizations could result in the incorrect extraction of the message.
@@ III-B Ambiguity Pool
where tokens $w_{i}$ and $w_{j}$ are in the candidate pool $V^{(t)}$. In the following algorithm description, we use $\textsc{prefix}\left(w_{i},w_{j}\right)=1$ to represent the presence of a prefix relationship between two tokens.
@@ III-E Extraction
III-F Proof of Security
@@
$\left|\text{Pr}\left[\mathcal{A}_{\mathcal{D}}({x_{s}})=1\right]-\text{Pr}\left[\mathcal{A}_{\mathcal{D}}({x_{c}})=1\right]\right|<\text{negl}\left(k\right).$ (12)
where $x_{s}$ is the disambiguating steganographic stegotext and $x_{c}$ is the normally generated covertext, $k$ is the security parameter of the shared key $K$ (usually the length of $K$), and $\text{negl}\left(k\right)$ is a negligible function concerning $k$.
@@
In our method, a token to be added to the stego sequence undergoes two sampling processes. The first sampling occurs within the distribution $P_{amb}$ formed after grouping by the ambiguity pool, controlled by the steganographic embedding algorithm and driven by the secret message. The result is a certain ambiguous pool $v_{amb}$, consisting of one or more ambiguous tokens. The second sampling occurs within $v_{amb}$, driven by pseudorandom numbers $r$ generated by a CSPRNG. For simplicity, we use $S(r,P)$ to represent sampling from $P$ using $r$, and $\mathcal{E}(m,P)$ to represent the steganography algorithm $\textsc{Encode}_{P}(\cdot,m)$, where $m$ is the secret message encrypted using a cryptographic algorithm. Therefore, the adversary’s judgement on $w_{s}$ can be expressed using the law of total probability as
@@
where $w_{s}=S(r,p_{amb}^{w_{s}})$, $v_{amb}^{w_{s}}=\mathcal{E}(m,P_{amb})$. At the same time, as the probabilities of each token remain unchanged before and after being grouped, a single normal generation of samples is also equivalent to two separate sampling processes using random numbers. The adversary’s judgment on $w_{c}$ can also be expressed as
@@
We also have $v_{amb}^{w_{c}}=S(r_{1},P_{amb})$ and $w_{c}=S(r_{2},p_{amb})$, where $r_{1}$ and $r_{2}$ are independent random variables corresponding to the equivalent two-stage sampling, each uniformly distributed, i.e. $r_{1},r_{2}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}U\left[0,1\right)$.
Upon comparison, it can be observed that to satisfy Eq. (14), it must be the condition that: the adversary can distinguish between $\mathcal{E}(m,P_{amb})$ and $S(r_{1},P_{amb})$ in polynomial time, or the adversary can distinguish between $S(r,p_{amb}^{w_{s}})$ and $S(r_{2},p_{amb})$ in polynomial time. Note that the provable security of the steganography algorithm we employ guarantees that the PPT adversary cannot have a negligible advantage $\delta^{{}^{\prime}}$ in distinguishing between $\mathcal{E}(m,P_{amb})$ and $S(r_{1},P_{amb})$, that is
$\left|\text{Pr}\left[\mathcal{A}_{P_{c}}({v_{amb}^{w_{c}}})=1\right]-\text{Pr}\left[\mathcal{A}_{P_{c}}({v_{amb}^{w_{c}}})=1\right]\right|<\text{negl}\left(k\right).$ (17)
@@ IV Experiments
Discop 16 3.27 0 0 1.30 0.84 480.10 3.68E-02 1.81
32 3.50 0 0 1.44 0.85 479.57 3.32E-02 2.80
64 3.88 0 0 1.58 0.85 480.84 3.04E-02 2.74
128 4.20 0 0 1.70 0.87 484.49 2.84E-02 2.50
256 4.55 0 0 1.81 0.88 487.83 2.69E-02 2.67
Discop + Basic 16 12.13 30.25 98.63 1.58 0.86 477.83 3.02E-02 0
32 18.75 37.34 98.64 1.93 0.86 483.34 2.51E-02 0
64 29.67 44.02 98.63 2.38 0.86 488.30 2.05E-02 0
128 42.26 49.10 98.63 2.80 0.85 508.60 1.82E-02 0
256 63.73 53.17 98.63 3.18 0.87 580.46 1.83E-02 0
MWIS 16 2.64 2.72 64.64 1.01 0.77 489.93 4.87E-02 0
32 2.83 2.94 71.56 1.12 0.76 500.53 4.47E-02 0
64 3.04 3.41 72.96 1.22 0.77 525.99 4.31E-02 0
128 3.22 3.58 68.56 1.29 0.76 986.66 7.66E-02 0
256 3.52 3.79 72.34 1.38 0.77 3765.48 2.73E-01 0
SyncPool 16 3.19 0 0 1.00 0.66 484.14 4.84E-02 0
32 3.74 0 0 1.03 0.59 487.56 4.75E-02 0
64 3.88 0 0 0.85 0.46 488.99 5.74E-02 0
128 4.50 0 0 0.63 0.31 488.00 7.73E-02 0
256 4.96 0 0 0.39 0.18 491.68 1.27E-01 0
@@
Discop 16 8.16 0 0 2.03 0.88 693.96 3.42E-02 1.73
32 10.97 0 0 2.56 0.90 691.17 2.70E-02 4.36
64 13.96 0 0 2.98 0.91 690.69 2.31E-02 5.36
128 20.12 0 0 3.60 0.93 692.19 1.92E-02 5.32
256 26.17 0 0 3.99 0.94 1000.93 2.51E-02 6.30
Discop + Basic 16 12.96 25.54 98.63 1.97 0.87 841.25 4.28E-02 0
32 20.22 30.34 98.63 2.48 0.86 867.03 3.50E-02 0
64 34.87 36.12 98.63 3.03 0.89 906.46 2.99E-02 0
128 45.63 38.36 98.59 3.32 0.87 950.55 2.86E-02 0
256 67.07 40.91 98.62 3.68 0.87 1056.15 2.87E-02 0
MWIS 16 6.63 6.89 62.76 1.60 0.75 705.43 4.42E-02 0
32 8.71 8.20 61.90 2.05 0.77 712.56 3.48E-02 0
64 12.15 9.87 64.48 2.55 0.78 723.59 2.84E-02 0
128 14.82 10.85 65.60 2.86 0.78 826.53 2.89E-02 0
256 19.50 11.92 67.10 3.23 0.78 2131.85 6.60E-02 0
SyncPool 16 7.49 0 0 1.57 0.72 695.04 4.42E-02 0
32 10.93 0 0 2.08 0.74 692.46 3.33E-02 0
64 14.40 0 0 2.46 0.74 693.70 2.82E-02 0
128 19.16 0 0 2.78 0.73 695.98 2.50E-02 0
256 23.57 0 0 3.05 0.73 700.90 2.30E-02 0
@@ IV-A Setup
In the experiments, we choose to conduct steganography experiments on English and Chinese, two of the most widely used languages worldwide. Meanwhile, unlike English, Chinese is a scriptio continua without spaces between characters. Researchers believe that segmentation ambiguity has a more significant impact on scriptio continua [26]. In our experiments, we will also explore this by comparing the results of experiments conducted in English and Chinese.
we deploy the pretrained models LLaMA2-7b  [^2] https://huggingface.co/meta-llama/Llama2-2-7b and Baichuan2-7b  [^3] https://huggingface.co/baichuan-inc/Baichuan2-7B-Base from Hugging Face as the basic generative model for English and Chinese, respectively. Both tokenizers are implemented based on subwords.
We introduce top-$k$ sampling to constrain the size of the initial candidate pool and the probability distribution. To better evaluate the performance in various scenarios, we set the truncation parameter $k=16,32,64,128,256$. For each $k$, we select $100$ pieces of text from the IMDb dataset [40]. We use the first three English sentences of each sample and their Chinese translations as the context for the LLaMA2 [38] and Baichuan2 [41] models to generate tasks in different languages, and generate the subsequent 100 tokens conditioned on this context.
@@ IV-B Metrics
We assess the performance of disambiguation algorithms primarily based on three aspects: the efficacy of disambiguation, its influence on the security of the steganographic method, and its impact on the efficiency of the steganographic method. We evaluate the impact by comparing the steganographic performance before and after the integration of the disambiguation algorithm using metrics for assessing steganographic methods.
@@ IV-B2 Security
In generative linguistic steganography, our goal is to ensure that the stegotext, where the message is embedded, is indistinguishable from normally generated text where no message is embedded. We use KL divergence to measure the distribution discrepancy between the original distribution provided by the language model during each token sampling period and the distribution used for steganographic sampling after disambiguation and steganographic methods.

### v2

@@ Abstract
Recent research in provably secure neural linguistic steganography has overlooked a crucial aspect: the sender must detokenize stegotexts to avoid raising suspicion from the eavesdropper. The segmentation ambiguity problem, which arises when using language models based on subwords, leads to occasional decoding failures in all neural language steganography implementations based on these models. Current solutions to this issue involve altering the probability distribution of candidate words, rendering them incompatible with provably secure steganography. We propose a novel secure disambiguation method named SyncPool, which effectively addresses the segmentation ambiguity problem. We group all tokens with prefix relationships in the candidate pool before the steganographic embedding algorithm runs to eliminate uncertainty among ambiguous tokens. To enable the receiver to synchronize the sampling process of the sender, a shared cryptographically-secure pseudorandom number generator (CSPRNG) is deployed to select a token from the ambiguity pool. SyncPool does not change the size of the candidate pool or the distribution of tokens and thus is applicable to provably secure language steganography methods. We provide theoretical proofs and experimentally demonstrate the applicability of our solution to various languages and models, showing its potential to significantly improve the reliability and security of neural linguistic steganography systems.
@@ II-A Steganography System and Steganographic Security
$\left|\text{Pr}\left[\mathcal{A}_{\mathcal{D}}^{\textsc{Encode}_{\mathcal{D}}\left(K,\cdot,\cdot\right)}=1\right]-\text{Pr}\left[\mathcal{A}_{\mathcal{D}}^{\mathcal{O}_{\mathcal{D}}\left(\cdot,\cdot\right)}=1\right]\right|<\text{negl}\left(\kappa\right),$ (4)
where $\mathcal{O}_{\mathcal{D}}\left(\cdot,\cdot\right)$ is an oracle that can randomly sample from the channel distribution $\mathcal{D}$, $\kappa$ is the security parameter of the shared key $K$ (usually the length of $K$), and $\text{negl}\left(\kappa\right)$ is a negligible function concerning $\kappa$.
@@ II-C Segmentation Ambiguity
The commonly used high-performance Transformer-based language models, represented by the GPT series [23, 24, 37] and the LLaMA series [35, 38], mostly utilize subword tokenization [27, 28, 29] to model text. Subword tokenization is a technique in which a word is split into subwords, and these subwords are known as tokens. This technique is used because a generative language model needs to maintain a large vocabulary and complex word structures. The concept behind this is that frequently occurring words should be in the vocabulary whereas rare words are split into frequent subwords. Each token is assigned an ID as a numerical representation of the subword  [^1] Since the numeric IDs corresponding to tokens differ across different models, in this paper, we directly use the subword mapped by the token to represent the token. In this paper, we use the underscore symbol “_” to indicate a word boundary, which is generally equivalent to a space in English.. For example, the word “_unwanted” might be split into “_un”, “want”, and “ed”. The stego generated by the steganographic encoding algorithm $\textsc{Encode}_{\mathcal{D}}$ is essentially a sequence composed of stego tokens. The sender must detokenize it using a tokenizer into a stegotext before transmission. The reason is obvious because even in today’s prevalent use of generative models and generated text, transmitting tokens directly over a public channel is a highly suspicious behavior. For example, if the sender generates two stego tokens mapping to subwords “_any” and “thing”, the sender needs to detokenize them into the text “anything” before sending it to Bob. However, the issue is that common words like “anything” often exist as independent tokens “_anything” in the model’s vocabulary as well. As a result, a single piece of text can correspond to two or even more different token representations. This phenomenon is referred to as segmentation ambiguity. To replicate Alice’s generation process and extract the message using $\textsc{Decode}_{\mathcal{D}}$, Bob must retokenize the received text into tokens. But since both “_any” and “_anything” exist in the candidate pool, Bob cannot determine which token the sender embedded the message into. Clearly, the differing tokenizations could result in the incorrect extraction of the message.
@@ III-B Ambiguity Pool
where tokens $w_{i}$ and $w_{j}$ are in the candidate pool $V^{(t)}$. $V^{(t)}$ represents the remaining candidate pool during the $t$-th time step of the generative model’s sampling process. All occurrences of $(t)$ in the following text have the same meaning, representing the time step applicable to the corresponding data. In the following algorithm description, we use $\textsc{prefix}\left(w_{i},w_{j}\right)=1$ to represent the presence of a prefix relationship between two tokens.
@@ III-E Extraction
III-F Complexity
Let $V$ denote the candidate pool and $|V|$ denote the size of the given pool. When SyncPool is used, the disambiguation algorithm performs an extra sorting operation and a traversal of all tokens in the candidate pool to compare common prefixes. The complexity of the sorting operation is $O(|V|\log(|V|))$, and the complexity for comparing common prefixes is $O(|V|)$. After this, the size of the candidate pool is reduced to $|V_{amb}|$. The complexity of steganographic embedding of the candidate pool changes accordingly, as determined by the steganographic algorithm used.
III-G Proof of Security
@@
$\left|\text{Pr}\left[\mathcal{A}_{\mathcal{D}}({x_{s}})=1\right]-\text{Pr}\left[\mathcal{A}_{\mathcal{D}}({x_{c}})=1\right]\right|<\text{negl}\left(\kappa\right).$ (12)
where $x_{s}$ is the disambiguating steganographic stegotext and $x_{c}$ is the normally generated covertext, $\kappa$ is the security parameter of the shared key $K$ (usually the length of $K$), and $\text{negl}\left(k\right)$ is a negligible function concerning $\kappa$.
@@
In our method, a token to be added to the stego sequence undergoes two sampling processes. The first sampling occurs within the distribution $P_{amb}$ formed after grouping by the ambiguity pool, controlled by the steganographic embedding algorithm and driven by the secret message. The result is a certain ambiguous pool $v_{amb}$, consisting of one or more ambiguous tokens. The second sampling occurs within $v_{amb}$, driven by pseudorandom numbers $r$ generated by a CSPRNG. For simplicity, we use $\mathcal{S}(r,P)$ to represent sampling from $P$ using $r$, and $\mathcal{E}(m,P)$ to represent the steganography algorithm $\textsc{Encode}_{P}(\cdot,m)$, where $m$ is the secret message encrypted using a cryptographic algorithm. Therefore, the adversary’s judgement on $w_{s}$ can be expressed using the law of total probability as
@@
where $w_{s}=\mathcal{S}(r,p_{amb}^{w_{s}})$, $v_{amb}^{w_{s}}=\mathcal{E}(m,P_{amb})$. At the same time, as the probabilities of each token remain unchanged before and after being grouped, a single normal generation of samples is also equivalent to two separate sampling processes using random numbers. The adversary’s judgment on $w_{c}$ can also be expressed as
@@
We also have $v_{amb}^{w_{c}}=\mathcal{S}(r_{1},P_{amb})$ and $w_{c}=\mathcal{S}(r_{2},p_{amb})$, where $r_{1}$ and $r_{2}$ are independent random variables corresponding to the equivalent two-stage sampling, each uniformly distributed, i.e. $r_{1},r_{2}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}U\left[0,1\right)$.
Upon comparison, it can be observed that to satisfy Eq. (14), it must be the condition that: the adversary can distinguish between $\mathcal{E}(m,P_{amb})$ and $\mathcal{S}(r_{1},P_{amb})$ in polynomial time, or the adversary can distinguish between $\mathcal{S}(r,p_{amb}^{w_{s}})$ and $\mathcal{S}(r_{2},p_{amb})$ in polynomial time. Note that the provable security of the steganography algorithm we employ guarantees that the PPT adversary cannot have a negligible advantage $\delta^{{}^{\prime}}$ in distinguishing between $\mathcal{E}(m,P_{amb})$ and $\mathcal{S}(r_{1},P_{amb})$, that is
$\left|\text{Pr}\left[\mathcal{A}_{P_{c}}({v_{amb}^{w_{c}}})=1\right]-\text{Pr}\left[\mathcal{A}_{P_{c}}({v_{amb}^{w_{c}}})=1\right]\right|<\text{negl}\left(\kappa\right).$ (17)
@@ IV Experiments
Discop 16 3.27 0 0 1.30 0.84 1.96 1.50 1.81
32 3.50 0 0 1.44 0.85 2.25 1.56 2.80
64 3.88 0 0 1.58 0.85 3.24 2.05 2.74
128 4.20 0 0 1.70 0.87 4.31 2.53 2.50
256 4.55 0 0 1.81 0.88 7.88 4.35 2.67
Discop + Basic 16 12.13 30.25 98.63 1.58 0.86 4.23 2.60 0
32 18.75 37.34 98.64 1.93 0.86 5.73 2.74 0
64 29.67 44.02 98.63 2.38 0.86 10.39 4.29 0
128 42.26 49.10 98.63 2.80 0.85 25.39 8.92 0
256 63.73 53.17 98.63 3.18 0.87 79.60 24.56 0
MWIS 16 2.64 2.72 64.64 1.01 0.77 13.32 11.61 0
32 2.83 2.94 71.56 1.12 0.76 20.39 15.53 0
64 3.04 3.41 72.96 1.22 0.77 37.56 26.79 0
128 3.22 3.58 68.56 1.29 0.76 133.83 85.04 0
256 3.52 3.79 72.34 1.38 0.77 1182.07 757.59 0
SyncPool 16 3.19 0 0 1.00 0.66 3.90 3.85 0
32 3.74 0 0 1.03 0.59 4.53 4.60 0
64 3.88 0 0 0.85 0.46 5.37 6.15 0
128 4.50 0 0 0.63 0.31 6.33 10.25 0
256 4.96 0 0 0.39 0.18 8.88 22.33 0
@@
Discop 16 8.16 0 0 2.03 0.88 2.00 1.53 1.73
32 10.97 0 0 2.56 0.90 2.33 1.61 4.36
64 13.96 0 0 2.98 0.91 2.90 1.83 5.36
128 20.12 0 0 3.60 0.93 4.16 2.44 5.32
256 26.17 0 0 3.99 0.94 7.92 4.37 6.30
Discop + Basic 16 12.96 25.54 98.63 1.97 0.87 4.52 2.77 0
32 20.22 30.34 98.63 2.48 0.86 6.17 2.96 0
64 34.87 36.12 98.63 3.03 0.89 10.87 4.49 0
128 45.63 38.36 98.59 3.32 0.87 26.33 9.25 0
256 67.07 40.91 98.62 3.68 0.87 83.13 25.65 0
MWIS 16 6.63 6.89 62.76 1.60 0.75 13.17 11.48 0
32 8.71 8.20 61.90 2.05 0.77 19.86 15.13 0
64 12.15 9.87 64.48 2.55 0.78 36.10 25.75 0
128 14.82 10.85 65.60 2.86 0.78 131.73 83.70 0
256 19.50 11.92 67.10 3.23 0.78 1048.35 671.89 0
SyncPool 16 7.49 0 0 1.57 0.72 4.11 4.06 0
32 10.93 0 0 2.08 0.74 4.80 4.87 0
64 14.40 0 0 2.46 0.74 5.95 6.83 0
128 19.16 0 0 2.78 0.73 8.07 13.05 0
256 23.57 0 0 3.05 0.73 12.77 32.12 0
@@ IV-A Setup
In the experiments, we choose to conduct steganography experiments on English and Chinese, two of the most widely used languages worldwide. Meanwhile, unlike English, Chinese is a scriptio continua without spaces between characters. Researchers believe that segmentation ambiguity has a more significant impact on scriptio continua [26]. In our experiments, we will also explore this by comparing the results of experiments conducted in English and Chinese. We deploy the pretrained models LLaMA2-7b  [^2] https://huggingface.co/meta-llama/Llama-2-7b-hf and Baichuan2-7b  [^3] https://huggingface.co/baichuan-inc/Baichuan2-7B-Base from Hugging Face as the basic generative model for English and Chinese, respectively. Both tokenizers are implemented based on subwords.
To more intuitively assess how performance is affected by the number of elements in the candidate pool, we only use top-$k$ sampling to constrain the size of the initial candidate pool and the probability distribution. The reason for not changing the temperature is that it has no effect on the number of tokens. As for top-$p$, its effect is similar to that of top-$k$, but the size of the candidate pool may fluctuate significantly at each sampling step, which is not easy to be quantified intuitively. In the experiments, we set the truncation parameter $k=16,32,64,128,256$. For each $k$, we select $100$ pieces of text from the IMDb dataset [40]. We use the first three English sentences of each sample and their Chinese translations as the context for LLaMA2 [38] and Baichuan2 [41] in different languages, and generate the subsequent 100 tokens conditioned on this context.
@@ IV-B Metrics
We assess the performance of disambiguation algorithms primarily based on three aspects: the effectiveness of disambiguation, its influence on the security of the steganographic method, and its impact on the efficiency of the steganographic method. We evaluate the impact by comparing the steganographic performance before and after the integration of the disambiguation algorithm using metrics for assessing steganographic methods.
@@ IV-B2 Security
In generative linguistic steganography, our goal is to ensure that the stegotext, where the message is embedded, is indistinguishable from normally generated text where no message is embedded. We use KL divergence (KLD) to measure the distribution discrepancy between the original distribution provided by the language model during each token sampling period and the distribution used for steganographic sampling after disambiguation and steganographic methods.
@@
We also use average perplexity to measure the perceptual imperceptibility of steganographic texts.
  • Ave PPL: the average perplexity of generated texts.
The smaller the difference in average perplexity values between the steganographic texts generated before and after applying the disambiguation algorithm, the stronger the security of the disambiguation algorithm. Ave PPL can be calculated as:
Ave $\textbf{Ave PPL}=\frac{1}{M}\sum_{j=1}^{M}2^{-\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log_{2}{\text{Pr}_{j}\left[x_{i}|x_{<i}\right]}},$ (18)
where $M$ is the number of text sequences tested, $N$ is the length of the sequence, $\text{Pr}_{j}\left[x_{i}|x_{<i}\right]$ is the probability of generating the $i$-th token given the previous tokens of the $j$-th text.